These are all part of the same question if you cant answer them all dont asnwer at all. I will thumbsdown. thanksLet y denote the number of broken eggs in a randomly selected carton of one dozen eggs. Suppose that the probability distribuSuppose that fund-raisers at a university call recent graduates to request donations for campus outreach programs. They reporAirlines sometimes overbook flights. Suppose that for a plane with 100 seats, an airline takes 110 reservations. Define the v
Answer #1

#1)

a) Total of probabilities P(y) of all y's must be equal to 1

0.64+0.20+0.11+0.04 + P(4) = 1

0.99 +P(4) = 1

Therefore P(4) = 1 - 0.99

P(4) = 0.01

b) P(1) = 0.2

In the long run , the proportion of cartons that have exactly one broken egg will equal 0.20

c) P( y <= 2 ) = P(0) +P(1)+P(2) = 0.64+ 0.20 + 0.11

P( y <= 2 ) = 0.95

Interpretation :

If you check a large number of cartons , the proportion that will have at most two broken eggs will equal to 0.95

d) P( y < 2 ) = P(0) +(1) = 0.64+0.20 = 0.84

This probability is less than the probability in part c) because the event y=2 is now not included.

e) The carton contain total 12 eggs , so P( exactly 10 unbroken eggs) = P( 2 broken eggs )

P(y = 2 ) = 0.11

#2)

1000 P(x) 400 P(x) 0.4 0.3 0.25 0.05 10 250 Total 1000

You would expect roughly 400 of the graduates to donate nothing, roughly 300 to donate $10, roughly 250 to donate $25,and roughly 50 to donate $50. the frequencies would be close to but not exactly , these values.The four frequencies would add to 1000

b) Most common value of x in this population is $0 ( because its proportion is higher )

c)P( x >= 25 ) = P( 25 ) + P( 50 ) = 0.25 + 0.05

P( x >= 25 ) = 0.30

c) P( x > 0 ) = P( 10 ) + P( 25 ) + P( 50 ) = 0.30 + 0.25 + 0.05

P( x > 0 ) = 0.60

#3)

a) P( airline can accommodate everyone who shows up ) = P(95) +P(96)+P(97)+P(98)+P(99)+P(100 )

= 0.04+0.09+0.12+0.16+0.21+0.17 = 0.79

b) P( not all passengers can be accommodated ) = 1 - P( airline can accommodate everyone who shows up )

= 1 - 0.79 = 0.21

c) P (number 1 standby will be able to take the flight ) = P(95) +P(96)+P(97)+P(98)+P(99)

= 0.04+0.09+0.12+0.16+0.21

= 0.62

d) P( number 3 standby will be able to take the flight ) = P(95) +P(96)+P(97)

= 0.04+0.09+0.12

=0.25